OpenStudy (anonymous):

Find an equation in standard form for the hyperbola with vertices at (0, ±4) and asymptotes at y = ±1 divided by 3.x.

OpenStudy (anonymous):

@Phebe do you know how to this one by any chance?

OpenStudy (anonymous):

ick another one!

OpenStudy (anonymous):

I hate this class haah almost done though 4 more haha

OpenStudy (anonymous):



OpenStudy (phebe):

let me see hold on k

OpenStudy (phebe):

yes i do

OpenStudy (anonymous):

because of the vertices, you know it looks something like that that means the \(y^2\) term comes first and it also means the center is at \((0,0)\) so it will look something like \[\frac{y^2}{a^2}-\frac{x^2}{b^2}=1\]now we only need \(a^2\) and \(b^2\)

OpenStudy (phebe):

but its seems @satellite73 is helping you XD

OpenStudy (anonymous):

no no go ahead, i hate these too, but we are really only two steps from done go ahead an finish it

OpenStudy (phebe):

i hate these too thas y im given it to you @satellite73

OpenStudy (anonymous):

ahh help? anyone haha

OpenStudy (phebe):

ok i will

OpenStudy (phebe):

hold on k

OpenStudy (phebe):

idk this one sorry

OpenStudy (anonymous):

It's okay thank you c:

OpenStudy (anonymous):

@satellite73

OpenStudy (phebe):

np